thrice the age of felicia is four more than the age of asare in three years time the sum of their ages is 30. find their ages
Answer by ankor@dixie-net.com(22740)

let f = Felicia's age
let a = Asare's age
:
Write an equation for each statement
:
Thrice the age of Felicia is four more than the age of Asare.
3f = a + 4
In three years time the sum of their ages is 30.
(f+3) + (a+3) = 30
f + a + 6 = 30
f + a = 30 -6
f + a = 24
Arrange it as:
f = -a + 24
in the first equation replace f with (-a+4)
3(-a+24) = a + 4
-3a + 72 = a + 4
72 - 4 = a + 3a
68 = 4a
a = 68/4
a = 17 yrs is Asare's age
then
f = -17 + 24
f = 7 yrs Felicia's age
:
:
Confirm this solution in the 2nd statement
"in three years time the sum of their ages is 30"
7 + 3 + 17 + 3 = 30